OpenStudy (anonymous):

the sum of a number and twice its reciprocal is33/4. Find the number

OpenStudy (anonymous):

\[x + 2*x^-1 = 33/4\]

OpenStudy (anonymous):

If you want it all to be in the power type ^{-1}

OpenStudy (anonymous):

that should be x + 2* x^(-1) = 33/4

OpenStudy (anonymous):

OR write: \[x + \frac{2}{x} = \frac{33}{4} \] \[\iff \frac{x^2 + 2}{x} = \frac{33}{4} \]

OpenStudy (anonymous):

how do you do the fractions? i've been trying to do it but i cant

OpenStudy (anonymous):

\frac{a}{b} gives \[ \frac{a}{b} \]

OpenStudy (anonymous):

And the if anf only if implication is \iff ... if anyone cares :(

OpenStudy (anonymous):

\frac{2}{3}

OpenStudy (anonymous):

It has to be in the equation editor

OpenStudy (anonymous):

ehhh\[\frac{2}{2}\] ahhhhhh thanks :D \[\iff\] kewlll thanks a lot